Consider the following information.

Stock price $55/share Avg. common shares outstanding 850,000
Common dividends $765,000 Dividends per common share $0.90/share
Preferred dividends $238,800 Net income $15,485,000
2019 preferred stock $6,162,000 2018 preferred stock $5,299,000
2019 total stockholders’ equity $28,435,000 2018 total stockholders’ equity $25,483,000
Purchases of treasury stock $120,000

Required:

1. Calculate the stockholder payout ratios. (Note: Round answers to three decimal places.)

Dividend Yield %
Dividend Payout %
Total Payout %
Stock Repurchase Payout %

2. Calculate the stockholder profitability ratios. (Note: Round answers as follows: Return on Common Equity to three decimal places; Earnings per Share to two decimal places.)

2019 Return on Common Equity %
Earnings per Share $ per share
Answer #1

1)

Formula
Dividend yield Dividend per share /market price per share

.90/55

1.636%

Dividend payout ratio Total common stock Dividend /Earning available to common stockholders

765000/15246200

5.018%

Total payout ratio Total dividend paid to both stockholders /Net income

(765000+238800)/15485000

1003800/15485000

6.482%

Stock repurchase payout Purchase of treasury stock/market capitalization of common stock

120000/(850000*55)

120000/46750000

.257%

**

Earning available to common stockholders = Net income -preferred dividend

                               = 15485000 - 238800

                               = 15246200

b)Common stockholders equity :

Beginning = Total -preferred

              = 25483000-5299000

              = 20184000

Ending= 28435000-6162000

            =22273000

Average common stockholders equity= [20184000 +22273000]/2

                             = 42457000/2

                                = 21228500

Formula
Return on common equity Earning available to common stockholders /Average common stockholders equity

15246200/21228500

71.819%

Earning per share Earning available to common stockholders/average common shares outstanding

15246200/850000

=$ 17.94 per share